OpenStudy (anonymous):

Simplify The Expression. -5+i/2i

OpenStudy (anonymous):

\[\frac{-5+i}{2i}\]\[\frac{(-5+i) \times i }{2i \times i}\]\[\frac{-5i-1}{-2}\]\[\frac{5i+1}{2}\]

OpenStudy (anonymous):

Thank You c: